# Intermittency and non-universality of pair dispersion in isothermal compressible turbulence

### v2

@@ Appendix C Lagrangian averages and dynamic exponents
Figure 10: Moments of Lagrangian exit times. Log-log plots of $\left\langle\tau_{\rm D}^{-p}(R)\right\rangle$ and $\left\langle\tau_{\rm H}^{-p}(R)\right\rangle$ versus $R$ for $p=2$ (blue), $p=4$ (red) and $p=6$ (violet) in runs (a) S1, (b) S2, (c) C1 and (d) C2. The vertical dashed lines demarkates the region of power–law fitting.
Figure 11: Convergence of Lagrangian averages. The dynamic exponents calculated using half ($5$) the members of the ensemble, compared with the ones calculated using the whole ensemble ($M=10$), displayed in Fig. 3 of the main text. The excellent overlap between the exponents and their errorbars, in every case affirms the convergence of the Lagrangian averages.

### v3

@@ B.5 Exit times
Figure 9: Moments of Lagrangian exit times. Log-log plots of $\left\langle\tau_{\rm D}^{-p}(r)\right\rangle$ and $\left\langle\tau_{\rm H}^{-p}(r)\right\rangle$ versus $r$ for $p=2$ (blue), $p=4$ (red) and $p=6$ (violet) in runs (a) S1, (b) S2, (c) C1 and (d) C2. The vertical dashed lines demarcate the region of power-law fitting in each case across which we extract $\chi^{\rm H}_{\rm p}$ and $\chi^{\rm D}_{\rm p}$.
Figure 10: Dynamic exponents for doubling and halving times $\chi^{\rm D}_{\rm p}$ and $\chi^{\rm H}_{\rm p}$, respectively, as functions of the order $p$ which takes integer and fractional values up to $6$. The blue shaded region corresponds to the bridge relation, $p-\zeta_{p}$, derived from the multifractal model. The pink shaded region corresponds to the expression $p-\zeta_{p}^{\rm s}$.
B.6 The multifractal spectrum
Figure 11: Multifractal spectrum. Plots of (a) $D_{\rm H}(h)$ and (b) $D_{\rm D}(h)$ versus $h$ for all the runs. $D_{\rm H}(h)$ changes slightly as we change the type of external force and Ma; the dependence of $D_{\rm D}(h)$ on Ma is much stronger in case of the solenoidal external force than the irrotational one. For S2, C1 and C2, $D_{\rm D}(h)$ takes values close to unity but not S1, implying that 1D structures contribute to the scaling of the doubling times in the former cases.
Following the multifractal model we consider the velocity field to comprise an infinity of interwoven sets, $\mathcal{S}_{\rm h}$, each characterized by a scaling exponent, $h$, and fractal dimension, $D(h)$. For any point ${\bm{x}}\in\mathcal{S}_{\rm h}$, $\delta u_{\parallel}(R)\sim R^{h}$, in the limit $R/L_{\rm I}\to 0$. We use this model to calculate the dynamic exponents, $\chi_{\rm{p}}$. The characteristic time scale (here, exit-time), $\tau_{\bm{x}}(R)\sim R/[\delta u_{\parallel}(R)]\sim R^{1-h}$. Its moments of negative orders, p-p ($p>0$), scale as
$\left\langle\tau^{-p}(R)\right\rangle\sim\int dw(h)\left(\frac{R}{L_{\rm I}}\right)^{d-D(h)-p(1-h)}\sim R^{-\chi_{\rm{p}}}\/;$ (16a)
$\implies-\chi_{\rm{p}}=\inf_{h}\left[ph+d-p-D(h)\right]\/$ (16b)
$\text{and}\quad D(h)=\inf_{p}\left(ph+d-p+\chi_{\rm{p}}\right)\/.$ (16c)
The spatial dimensionality $d=2$, and $dw(h)$ gives the weights of the different exponents. For $R/L_{\rm I}\to 0$, the integral in  (16a) is evaluated by the method of steepest descents. Equation (16c) implies that $D(h)$ and $\chi_{\rm{p}}$ are Legendre transforms of each other. Physically, $D(h)$ reveals the dimensionality of structures that contribute to the scaling of $\delta u_{\parallel}(R)$ with the exponent $h$, unveiling the multifractal spectrum. Note that, even if $D(h)$ is not concave its Legendre transform $\chi_{\rm{p}}$ will be concave; however, the inverse transform returns not $D(h)$ but its concave hull [see, e.g., 7, section 8.5].
B.7 Error analysis
For equal-time structure functions we take $M=10$, well-separated-in-time, snapshots, from the stationary state of our DNS. From each snapshot we calculate the structure function  (15), choose a scaling range by eye estimation (this range is the same is all the snapshots) and fit the log-log plot of the data with a linear function to obtain the equal time scaling exponents. In the worst case we obtain close to a decade of clean scaling. The structure functions from one such snapshot is shown in Fig. (6). Thus we obtain $M$ independent estimates for each exponent. The mean of these $M$ values is our estimates of the equal time exponets and the standard deviation is the error.
For the scaling exponents obtained from the exit times, we introduce $2n$ tracer pairs, where $n=4096^{2}$, in the statistically stationary state and select only thoese whose initial and their final separation are within the inertial range. Thence we calculate the negative moment of the exit times and plot them, in log-log scale, as a function of $R$ and use linear fit to estimate the dynamic scaling exponents. We repeat this exercise for $M=20$ independent initial realization of the tracer pairs. For each exponent, its final estimate is its ensemble mean, and its error bars are defined by the standard deviation. These are the values plotted in Fig. (10).
This method of estimating parameters of a model and its errors from an ensemble based approach is often called BAGGing or Bootstrap AGGregation method, see e.g., Ref. [51] section 8.2.